Researchers have created any index they claim measures an adult male's propensity to commit an act of road rage. This index is calculated from the answers to a questionairre and the scores are normally distibuted amoung US adult males, with a mean of 100 and a standard deviation of 30.
Researchers assess a group of twenty adult male drivers convicted of committing road rage. The mean score for the group is 50.8.
What percentage of US adult males would be expected to obtain a score equal to or less than that of road-ragers?

Answer by stanbon(75887)

Researchers have created any index they claim measures an adult male's propensity to commit an act of road rage.
----------------------
This index is calculated from the answers to a questionairre and the scores are normally distibuted amoung US adult males, with a mean of 100 and a standard deviation of 30.
----------------------
Researchers assess a group of twenty adult male drivers convicted of committing road rage. The mean score for the group is 50.8.
----------------------
What percentage of US adult males would be expected to obtain a score equal to or less than that of road-ragers?
-----
x = zs+u
50.8 = z*30 + 100
30z = -49.2
z = -49.2/30 = -1.64
----
P(x <= 50.8) = P(z <= -1.64) = normalcdf(-100,-1.64) = 0.0505 = 5.05%
=======================================================================
Cheers,
Stan H.